Bilal, Charlene, Caroline, Iacopo, Florian, Stefano, Carsten arrive at the 30m. Full house.
Four shifts this time:
Afternoon (12:00-18:00 CET): Charlene and Caroline
Evening (18:00-00:00 CET): Bilal
Night (00:00-6:00 CET): Florian and Iacopo
Morning (6:00-12:00 CET): Carsten and Stefano
Last week NIKA2 was cooled down a bit earlier than normal, to allow doing a new frequency sweep at 1mm. Meanwhile, the pulse tube has been realigned by JLS, after notification by Martino who saw indications of a spike at 0.3 Hz in the power spectrum.
Jan 14, Tuesday
Maintenance and afternoon/evening shifts
- Laser used to check alignment between elevation axis and cryostat window. Moving in elevation, the laser rotates, and this may explain small offsets on the cryostat window. Small correction applied to M5 by SN. A test of shining the laser from the window to M2, and moving in elevation, was not made.
- Starting at 14:30 UT. High overcast of ice clouds.
- Pointing session with ~33 sources (scans 237-269). Pointing offsets look OK. JP fits model. Fit of Nasmyth offsets leads to a very slight improvement from 1.9" rms to 1.7" rms. Not yet implemented. Beam widths look OK.
- Secondary calibrators MWC349, NGC7027, CRL2688. Fluxes look OK.
- Beam map on Uranus at 50deg Elevation.
- Central pixel lost. Hole in the 2mm array. Re-do tuning. Re-do pointing, focus, 2 beam maps to create new RPPs/kidpars after 1mm sweep done last week (scans 292 and 293).
- tau225~0.14 and very stable. Forecast is good for the next two days. Ready to go. It would be helpful to have beam maps IMBfits be created separately from pointing and focus scans.
18:10 Finally, after 3.5 hours, starting with science: NIKA2SZ. New pointing model implemented in-between scans.
- Alternating NIKA2SZ and N2CLS COSMOS, with some observations of 128-19 and 126-19 too. Tau very stable, slowly degrading.
Jan 15, Wednesday
What a day!
- Continuing with N2CLS GOODS-North. Pointing, focus, tau stable. tau225=0.2 @ 05:30 UT.
8:00 UT: Juan P. informs us know that the pulse tube has been re-aligned. Santiago confirms via email and explains the details: a 0.3 Hz spike was found remotely by Martino (Jan 9th) and Juan Louis at the telescope re-aligned the pulse tube.
At 9:00 UT Martino writes an email and lets us know that: " the levels of the cryostat (i.e. its position wrt to the floor) are not well in line with the 'reference' values. This is due to the fact that the cryostat is in a 'manual' state: no checks are done on it and no pressure adjustments in the pressure of the NIKA2 pneumatic feet are done. The state should be changed to 'level attente'. But.. This will move a little bit the cryostat! (...) For the time being, I don't touch anything, but in my opinion it would be better to go back to the standard operatin mode (level attente)." Comment by SB: At the moment, the pointing is quite stable and the sky is good too. Yesterday the mirrors were re-aligned and a new pointing model was done. I guess that if the cryostat is moved (by setting the leveling to "automatic"), then a new re-alignment will be necessary and shall be done. And a new pointing model too. Is it necessary right now, or can we wait the end of the run?
- The operator (Ignazio) tries to start the cryostat monitor here but the connection does not work. Santiago will work it out later (he's now in a meeting). Yesterday it was on, as far as Ignazio remembers. Bilal confirms.
- After some interaction with Martino, Santiago, Miguel, we decide to postpone switching to the automatic leveling of the cryostat to next Tuesday (end of the run), unless big pointing problems arise.
- Meanwhile the connection to the cryostat monitor is re-established here too.
at 10:45 UT, synchronization problems between Boxes A and M/N/T (scan 81 of today) are reported by Nicolas via Juan and Martino. He used the IDL RTA. We cannot check with the PIIC monitor right now, because the RPPS for the new sweep are currently on the make by Stefano (who had some interruptions meanwhile, sorry). We reboot the electronic boxes after scan 92.
- After the reboot, we observe project 143-19
At 12:10 UT connection was lost to the cryostat control machine. Valves closed automatically. The problem was realized around 12:50 UT. NIKA2 warmed-up to 500mK, when control was regained. Observations are stopped until operating temperature is reached again and is stablized.
- At ~13:45 UT, and 0.178 K, the tuning is again effective and we can start to observe again. The temperature will decrease further, down to ~0.17. THe observer is asked if she is willing to continue the observations in these conditions. The observations of project 143-19 are resumed when the mixing chamber reaches a temperature of 0.171 K.
- The amps and the temp regulation has been switched on at ~14:45 CET by Dave
- In total 4 sources of 143-19 (two before and two after the temperature "break")
- After some struggling to find the proper focus (it's afternoon after all), we switch to project 160-16. Two more scans on ngc891 (tau225~0.15 and stable).
When scans are long, the make-imbfits has delays. It would be nice to produce imb-fits of large scans in parallel, without slowing down the short ones (e.g. pointing and focus taken between or after long scans).
- We observe projects 160-16, 128-19, 163-19.
- Continuing with 163-19 and LPSZ.
Array #3 stops working after midnight UT.
- All attempts to get it working again fail: retune; restart the electronics; hard reboot and empty the capacitors.
- the PIIC monitor stops accepting the Ar2 data. Continuing with the IDL RTA.
- About 5 hours lost.
- The AOD decides to continue observations with Ar2 and Ar1 only, in order to still use the nice sky conditions we have.
- Observing 219-19 from the account nikaw-19.
Jan 16, Thursday
Today wasn't bad either...
- How to account observing time (using Ar2 and Ar1 only), shall be clarified later.
- Ar3 still "dead"
- Trying to investigate why the PIIC monitor does not accept the Ar2 data, with no success
- Using the IDL RTA; for pointing we use Ar2 only, as usual
- Observing 219-19. Stopping 219-19 at 05:40, as it has not been in the NIKA2 pool.
- 05:45 UT. We switch to project 123-19 (under account 123-19): observing source L134. tau225~0.22.
- 07:00 UT: tau225~0.2. Blue sky. No clouds. No wind.
07:00 UT: Robert intervenes on PIIC. Not clear what the problem was; killing the monitors & restarting helped. Using the RPPs of the individual maps produced yesterday by SB for the new ar1&3 sweep, RZ calculated the merged RPPs and installed them. The PIIC monitor processes 1mm data again.
08:20 UT we interrupt the observations for a short time, to recover Ar3 with the help of Santiago and Ignazio. 08:28UT: Synthesizer 3 switched off and on again, following instructions of SN. Trying to re-tune Array 3. Re-boot the acquisition and re-tune. Array 3 works again!!! it's 08:33 UT.
- starting from scan 51 of today, Ar3 is back.
- we are now also using the PIIC monitor again for pointing and focus.
- we continue observing project 123-19
- since we've not had enough problems yet, scan 53 (2nd scan of a focus sequence) has synchronization problems. This issue is reported both by the PIIC monitor and the IDL RTA. This does not hinder the focus determination.
- when having large scans, the make-imbf is really delaying the processing. For example when a pointing is taken after a long scan, we get to wait up to ~5 minutes before the IMBF are ready and the PIIC monitor can process them. This is even more severe at 1mm
So far, the morning shift had their share of problems (yesterday/today); the night shift had their share (last night); the afternoon shift had their share (yesterday); Bilal had his share (during the night)... I think we're done... from now on everything can proceed smoothly... uhm.. wait.. when is the snow storm coming???
- We take over and complete one more loop on L134 (proj. 123-19). In total, 6 loops were observed today (3 of them with Ar 1 and 2 only)
- Project 122-16 (GASTON LP) follows; only 22 deg from the sun
~14:30 UT Anomalous refraction: trying to focus and studying beam. Note by SB: the previous scans taken "close" to the sun might also play a role.
~15:30 UT we try project 219-19; the beam is still unstable.
- We observe projects 217-19, 128-19 and 163-19 in increasing opacities. The start of the evening was clear, but the beam was still unstable.
- We observe project 143-19.
- Switched to EMIR 3mm backup project at 3:30UT: 205-19, when tau225 went above ~0.4.
Jan 17, Friday
The Sun's show
- Continuing with EMIR. tau225~0.6. mrtcal gives 12mm pwv.
- 07:10UT tau225 rapidly improving. Now 0.3. After a change of EMIR tuning, pointing, focussing, reconsider.
- 07:40UT Changing to NIKA2. retuning ok. tau225~0.17.
Observing project 123-19 (source L134-1). tau225 all the way down to 0.14 . tau225~0.17 at 09:37UT very stable.
10:20UT Skydip (tau225~0.17)
- 160-16: NGC891: 4 more scans. NGC6946: 2 scans (1h). 18:30UT tau225~0.26 slowly rising.
- 218-19 20:30UT tau225~0.27
- the sky is degrading;
- we observe 212-19, which accepts poorer sky conditions
Jan 18, Saturday
- 05:00 UT At the beginning of the shift, we switch to EMIR; tau225=0.47
- ~06:30 UT we stop because the sky conditions are poor even at 3 mm. We wait half an hour and see how it evolves. After half an hour, the sky is too bad to continue. 07:30UT telescope put in stow position.
Stopped the remainder of the day.
Jan 19, Sunday
Stopped the full day. I think the run's over. We're deep into the snow storm...
Jan 20, Monday
Stopped the full day. High wind and some snowfall. No chance to use the platform.